f(x)=2x+5,g(x)=x^2-10 and h(x)=3x-8, then what is g(f(5)) and f(g(-n))
Answer by actuary(112)

f(5) = 2*10+5 = 25
So g(f(5)) = g(25) = 25^2-10 = 625-10 = 615
g(-n) = (-n)^2-10=n^2-10
So f(g(-n)) = f(n^2-10)= 2*(n^2-10) + 5 = 2*n^2-20+5 = 2*n^2-15